# Detecting AI-Generated Images via Diffusion Snap-Back Reconstruction: A Forensic Approach

### v1

@@ Abstract
The rapid rise of generative diffusion models has made distinguishing authentic visual content from synthetic imagery increasingly challenging. Traditional deepfake detection methods, which rely on frequency or pixel-level artifacts, fail against modern text-to-image systems such as Stable Diffusion and DALL-E that produce photorealistic and artifact-free results. This paper introduces a diffusion-based forensic framework that leverages multi-strength image reconstruction dynamics—termed diffusion snap-back—to identify AI-generated images. By analysing how reconstruction metrics (LPIPS, SSIM, and PSNR) evolve across varying noise strengths, we extract interpretable manifold-based features that differentiate real and synthetic images. Evaluated on a balanced dataset of 4,000 images, our approach achieves 0.993 AUROC under cross-validation and remains robust to common distortions such as compression and noise. Despite using limited data and a single diffusion backbone (Stable Diffusion v1.5), the proposed method demonstrates strong generalization and interpretability, offering a foundation for scalable, model-agnostic synthetic media forensics.
@@ I Introduction
In recent years, the widespread emergence of large-scale generative models has reshaped how visual media is produced and consumed. State-of-the-art diffusion models (such as Stable Diffusion and Midjourney), transformer-based architectures (such as DALL-E and Imagen), and their latent diffusion variants are now capable of producing highly realistic images across a vast range of domains—spanning landscapes, portraits, objects, and abstract scenes [1, 2]. At the same time, today’s social media ecosystem is saturated with synthetic imagery and video, such as tools like Google Veo 3 can generate video clips that are nearly indistinguishable from real footage, raising serious concerns about the erosion of trust in digital content [3].
This shift is particularly consequential for older or less digitally-savvy populations, who may struggle to distinguish real from fake content[4]. In Bangladesh, AI-generated and manipulated images and videos are increasingly used for political promotion, defamation of parties, and the spread of misinformation[5]. As visual content becomes easier and cheaper to produce and share, the challenge of verifying authenticity grows more urgent.
Unlike deepfakes, which typically manipulate existing visual content, modern diffusion-based systems can generate entirely novel and photorealistic scenes from textual prompts[6]. This fundamental shift renders earlier detection approaches—such as frequency-domain analysis, CNN-based artefact classification, or heuristic methods exploiting facial or eye-gaze inconsistencies—largely ineffective beyond face-centric data. Moreover, the statistical characteristics of diffusion-generated images differ substantially from those produced by generative adversarial networks (GANs), making the search for robust, domain-agnostic forensic indicators an urgent research priority[2].
To bridge this growing gap, our study explores how diffusion models behave when reconstructing real versus AI-generated content. Diffusion models, by design, learn to represent the distribution of synthetic images within a complex data manifold, whereas authentic, human-captured images usually exist outside that space. When these models are used for image-to-image reconstruction under varying noise levels $\mathcal{S}=\{0.15,\,0.30,\,0.60,\,0.90\}$, the difference becomes visible. Real images tend to lose perceptual quality abruptly as noise increases, largely because their formation process is complex, diverse, and influenced by countless natural and environmental factors that diffusion models cannot perfectly replicate after a certain period. But AI-generated images degrade more smoothly, maintaining structural and semantic consistency since they originate directly from the model’s learned manifold. This clear difference in reconstruction behavior forms the foundation of our approach to identifying synthetic imagery.
Looking ahead, such forensic techniques could support the development of a public verification platform—an accessible online space where anyone, from journalists to everyday social media users, could upload an image or video to check whether it was likely generated by AI. A system like this would be especially valuable in today’s misinformation-prone environments, including regions such as Bangladesh, where AI-generated visuals are increasingly used in political propaganda and public manipulation, and where digital literacy levels vary widely. The key contributions of this work are summarized below:
@@
    Diffusion-based forensic framework: We introduce a multi-strength img2img reconstruction method that extracts both point-wise (LPIPS, SSIM, PSNR) and curve-level (AUC-LPIPS, $\Delta_{LP}$, knee-step) features to distinguish real and AI-generated images.
@@
    Comprehensive evaluation: The method is validated on a balanced dataset of 4,000 authentic and synthetic images with stratified cross-validation, ablation studies, and robustness testing across common distortions such as compression, blur, and noise.
@@
    Interpretability and generalization: The proposed approach provides an interpretable, model-agnostic forensic signal directly tied to diffusion reconstruction behavior, enabling consistent performance across diverse visual domains.
@@ II-A Synthetic Image Detection
Early efforts in detecting synthetic media largely targeted images generated by GAN-based architectures, especially facial imagery. Pioneering works used frequency-domain anomalies or trained CNN classifiers on GAN-artifact cues [7, 8]. While these methods achieved promising performance in constrained settings, they often fail to generalise beyond face-centric datasets. Recent surveys observe that many GAN-era detectors struggle when applied to newer generative methods or non-face domains [9, 10]. In particular, the emergence of diffusion-based generators has drastically reduced visible frequency artifacts and introduced smoother textures with more plausible physical consistency, thereby undermining the assumptions of legacy detectors [11]. Moreover, many detection frameworks are trained on single generation methods and lack robustness to compression, blur, or unseen model distributions.
@@ II-B Diffusion Model Analysis and Forensics
Diffusion models, grounded in reverse Markov chains or stochastic differential equations, have become dominant in high-fidelity image generation [12, 13]. A handful of recent works explore their properties: for example, one study shows that pre-trained diffusion models themselves can act as implicit detectors of synthetic imagery via strategic sampling [14]. Another recent direction examines reconstruction error bounds for real versus generated images in the latent space of diffusion decoders [15]. However, despite these promising signals, the forensic community still lacks systematic frameworks that exploit diffusion-model behaviour (e.g., manifold membership, reconstruction dynamics) specifically for synthetic detection.
@@ II-C Manifold-Based Forensics and Feature Design
Traditional image forensics frequently relies on pixel- or frequency-level cues (such as camera sensor noise, JPEG artifacts) [16]. More recently, manifold-based reasoning has emerged: the idea that generative models implicitly learn a data manifold, and that synthetic images lie on or near it whereas real images lie off manifold [17, 18]. A few methods attempt to exploit this by modelling reconstruction or projection error into latent spaces. Yet many of these approaches either rely on costly inversion of generative models, or assume that the same generation model is used for detection. As a result, generalisation across unseen models and content types remains limited.
II-D Research Gap
While the literature has made considerable progress, especially in GAN-based detection and the analysis of diffusion models, several key gaps remain. First, most detectors do not generalise well beyond specific models or domains (e.g., faces) and are vulnerable to real-world distortions (compression, blur). Second, the majority of forensic methods still rely on handcrafted image‐level features rather than exploiting the dynamic reconstruction behaviour of diffusion models. Third, there is minimal work that treats diffusion models themselves as forensic sensors for general content types (objects, scenes, landscapes). Therefore, a compelling research direction is to design model-agnostic, dynamically‐informed features based on diffusion model reconstruction trajectories that generalise across content domains and generation architectures.
[…]
@@ III Methodology
Our detection framework relies on analysing how diffusion models reconstruct images under varying noise levels—a process we term diffusion snap-back. The methodology integrates feature extraction, classification, and robustness evaluation, implemented in PyTorch and scikit-learn with Stable Diffusion v1.5 as the generative backbone.
@@ III-A Dataset
We used the AI vs. Human-Generated Images Dataset by Alessandra Sala  [19], as the primary dataset for our experiments. The dataset contains equal proportions of authentic (human-captured) and AI-generated images across multiple categories such as portraits, objects, and scenes. Each image is labeled as 0 for human-generated content and 1 for AI-generated content, as illustrated in Table I. Representative samples from both classes are shown in Figure 2, where the top row depicts real images and the bottom row shows AI-generated counterparts.. Although primarily face-centric, the dataset encompasses diverse visual content such as portraits, landscapes, and objects, allowing the method to generalize across multiple domains. Authentic samples include photographs and original artwork, while synthetic samples are produced via Stable Diffusion v1.5.
All images are resized to $512\times 512$ RGB. Diffusion reconstruction employs Stable Diffusion v1.5 with the DDIM scheduler ($T=50$ inference steps), and reconstruction strengths $\mathcal{S}=\{0.15,0.30,0.60,0.90\}$ under a guidance scale $w=1.0$. All experiments are conducted on an NVIDIA RTX 4090 GPU using torch.float16 precision with optimization strategies such as attention slicing, VAE slicing, and CPU offload.
@@ III-C Classification Pipeline
The extracted features were used to train a lightweight logistic regression classifier with $\ell_{2}$ regularization. Prior to training, missing values were imputed using the median and features were standardized. Stratified five-fold cross-validation was conducted to assess generalization, yielding mean AUROC and AUPRC values. The optimal decision threshold $\theta^{*}$ was determined via Youden’s J-statistic, balancing sensitivity and specificity. To benchmark performance, a pixel-level baseline using 32×32 flattened image vectors achieved only 0.525 AUROC, underscoring the effectiveness of the proposed manifold features. On the full feature set, the snap-back model reached 0.993 cross-validation AUROC and 0.990 on a 35% holdout test split.
III-D Robustness Evaluation:
To evaluate real-world applicability, six augmentation conditions were applied to a balanced subset of 24 images (12 real, 12 AI-generated): raw, JPEG-60, WebP-60, Gaussian blur (radius 1.2), additive Gaussian noise ($\sigma=6.0$), and simulated screenshot resampling (downscale to 320 px and upsample). For each augmented variant, snap-back features were recomputed and AUROC was measured. Results showed high resilience under compression (83–87% AUROC) and moderate degradation under blur or screenshot perturbations, indicating robustness to common online distortions.
Overall, this methodology operationalizes diffusion reconstruction dynamics into a compact and interpretable feature space. By combining multi-strength perceptual metrics with statistical learning, it transforms diffusion models themselves into forensic sensors capable of differentiating authentic and AI-generated images.
@@ IV-A Primary Results
Our diffusion snap-back feature pipeline achieves an impressive AUROC of 0.993 on 5-fold cross-validation, and maintains 0.990 on a 35% holdout test set. The optimal decision threshold determined via Youden’s J-statistic is $\theta^{*}=0.914$, yielding balanced sensitivity and specificity.
The overall performance metrics are visualized in Fig. 5, which shows the ROC curve, calibration reliability plot, and confusion matrix at $\theta^{*}$. The model demonstrates high discriminative power and near-perfect calibration, with minimal false positives or negatives. The classifier demonstrates both high discriminative power and near-perfect calibration, with only minor misclassifications between human and AI-generated categories.
@@ IV-B Ablation Study
Ablation analysis identifies knee-step (the threshold at which SSIM drops below 0.8) as the single most discriminative feature. Combined with LPIPS at higher diffusion strengths and curve-level AUC summaries, the model approaches full-feature accuracy with minimal redundancy.
@@ IV-G Why Snap-Back Works
Diffusion models learn a denoising function $\nabla\log p_{\theta}(x)$ that approximates score-matching on the underlying data manifold. Real images, drawn from natural or human-created distributions, generally lie outside this learned manifold because they represent the full complexity and variability of the real world. In contrast, AI-generated images produced by models such as Stable Diffusion exist directly on, or very close to, the manifold that the model has learned during training.
When noise is injected and the reconstruction process begins, this distinction becomes evident. For real images, the addition of high noise levels drives the sample farther from its original representation, and the diffusion model—having never learned to denoise toward such off-manifold regions—struggles to recover the true structure. This results in higher LPIPS values and a sharp decline in SSIM. Conversely, AI-generated images remain within the learned manifold even after noise perturbation, allowing the model to denoise smoothly back to a plausible image representation, leading to lower LPIPS and gradual SSIM decay.
This difference in reconstruction behavior effectively encodes manifold membership and is largely model-agnostic: any diffusion model trained on natural image distributions is expected to display similar dynamics when processing in-distribution (synthetic) versus out-of-distribution (real) inputs. Consequently, the snap-back effect serves as a universal and interpretable signal for differentiating authentic from AI-generated content.
@@ V Conclusion
In this study, we introduced a diffusion-based forensic framework that distinguishes real images from AI-generated ones by observing how they behave during multi-strength image reconstruction. Even with a relatively small dataset of 4,000 images, our method achieved strong results—showing that diffusion reconstruction patterns can serve as reliable indicators of synthetic content. The findings highlight that real images tend to degrade abruptly under increasing noise, while AI-generated ones degrade more smoothly, reflecting their alignment with the learned diffusion manifold. However, this work also comes with several limitations. Our experiments were conducted using Stable Diffusion v1.5 as the core model, and we have not yet tested the framework on other diffusion systems such as SDXL, DALL-E 3, or Midjourney. The utilized dataset size is limited, and we lack access to large-scale, high-quality data that could improve the robustness and generalization of our results. Due to hardware constraints, we were also unable to train a diffusion model from scratch that could provide deeper insights into how synthetic and real images occupy different regions of latent space.
In future, we aim to extend this research to larger and more diverse datasets, incorporating new diffusion architectures and higher-quality AI-generated imagery. With better GPU resources, we plan to train a custom diffusion model from scratch to explore the latent manifold more directly. Beyond static images, we also hope to adapt this approach for videos, where temporal consistency could reveal even stronger forensic cues. Ultimately, our goal is to contribute to the growing need for trustworthy AI and reliable detection of synthetic media in an increasingly visual digital world.

### v2

@@ Abstract
The rapid advancement of generative image models has transformed digital media to the point where AI-generated images can no longer be reliably distinguished from authentic photographs by human observers or many conventional detection methods. Modern text-to-image systems such as Stable Diffusion and DALL-E can now generate images so realistic that they often appear completely natural, leaving little to no visible artifacts for traditional deepfake detectors to rely on. This challenge has practical consequences for misinformation control, institutional identity verification, and digital trust in political and legal contexts. Instead of searching for hidden pixel-level traces, we take a different approach: we observe how an image responds when it is gently disturbed and reconstructed by a diffusion model. We call this behavior diffusion snap-back. By tracking how perceptual similarity measures (LPIPS, SSIM, and PSNR) change across different reconstruction strengths, we capture compact and interpretable signals that reveal how closely an image aligns with the diffusion model’s learned denoising behavior. Evaluated on a balanced dataset of 4,000 human and AI-generated images, the proposed method achieves an AUROC of 0.993 under stratified five-fold cross-validation and 0.990 on a holdout split using only logistic regression. Initial robustness tests show that the method remains stable under common real-world distortions such as image compression and added noise. Although our experiments were conducted using a single diffusion backbone, the results indicate that reconstruction behavior can serve as a reliable and scalable foundation for synthetic media detection as generative models continue to grow more realistic.
Accepted at the 2026 IEEE 2nd International Conference on Quantum Photonics, Artificial Intelligence and Networking (QPAIN 2026). Copyright IEEE.
@@ I Introduction
Large-scale generative models have transformed how visual content is created and consumed. Modern diffusion systems such as Stable Diffusion, Midjourney, DALL-E and Gemini 3.1 Flash Image can synthesize highly realistic images across diverse domains [2, 18]. Recent multimodal tools capable of producing near-photorealistic video further blur the boundary between authentic and synthetic media [4]. As visual artifacts disappear, distinguishing real from generated content becomes increasingly difficult.
This technological shift carries serious societal and institutional risks. AI-generated visuals are now widely used in misinformation campaigns, political propaganda, and reputational attacks [6]. Older or less digitally literate populations may be particularly vulnerable [1]. Beyond public misinformation, synthetic imagery also raises concerns in everyday institutional settings. University or academic admission systems that rely on uploaded photographs may be vulnerable to impersonation or proxy examination schemes, where a synthetic or manipulated image is submitted online and a different individual later appears for the exam. Fabricated visuals can also influence criminal investigations or legal disputes, and manipulated images may fuel rumors in highly polarized political environments. These challenges highlight the need for scalable and reliable authenticity verification tools.
Unlike traditional deepfakes, diffusion-based systems generate entirely new scenes from learned distributions [11]. Consequently, many legacy detection methods—such as frequency-based artifact analysis or GAN-specific fingerprinting—fail to generalize. Diffusion-generated images often lack obvious statistical irregularities, motivating a shift from static artifact inspection to behavior-based analysis.
In this work, we examine how images respond when reconstructed by a diffusion model under controlled noise strengths $\mathcal{S}={0.15,0.30,0.60,0.90}$. Diffusion models refine noisy inputs through iterative denoising, and the way an image reconstruct reveals how closely it aligns with the model’s learned prior. We observe that AI-generated images typically degrade more smoothly as reconstruction strength increases, whereas authentic photographs often show sharper perceptual changes at higher noise levels. Rather than claiming strict separation, we interpret this difference as varying degrees of alignment with the denoising prior. We term this reconstruction behavior diffusion snap-back, which forms the basis of our forensic framework. The contributions of this work are as follows:
@@
    We introduce a reconstruction-dynamics detection framework that treats a pre-trained diffusion img2img pipeline as a forensic probe rather than relying on pixel-level artifacts.
@@
    We design a compact 15-dimensional feature representation combining multi-strength LPIPS, SSIM, and PSNR metrics with trajectory descriptors (AUC-LPIPS, $\Delta_{LP}$, knee-step).
@@
    We establish a lightweight classification pipeline based on logistic regression and stratified cross-validation to operationalize diffusion snap-back features for synthetic image detection.
  4. 4.
    We conduct ablation, correlation, and robustness analyses to assess feature contribution, interpretability, and stability under common distortions.
@@ II-A Synthetic Image Detection
Early efforts in synthetic media detection primarily focused on GAN-generated images, particularly facial content. Initial approaches leveraged frequency-domain anomalies or trained CNN classifiers to detect GAN-specific artifacts [10, 8]. Although effective in constrained settings, these methods often fail to generalize beyond face-centric datasets. Recent surveys report that many GAN-era detectors struggle when applied to newer generative models or broader visual domains [7, 12]. The rise of diffusion-based generators has further reduced visible frequency artifacts and introduced smoother textures with stronger physical consistency, weakening assumptions underlying legacy detectors [5]. Additionally, many detection frameworks are trained on specific generation methods and exhibit limited robustness to compression, blur, or unseen model distributions.
@@ II-B Diffusion Model Analysis and Forensics
Diffusion models, grounded in reverse Markov chains and stochastic differential equations, have become the dominant paradigm for high-fidelity image synthesis [17, 3]. Recent studies explore their analytical properties: some demonstrate that pre-trained diffusion models can act as implicit detectors of synthetic imagery via strategic sampling [16], while others investigate reconstruction error behavior in the latent space of diffusion decoders [9]. Despite these promising directions, systematic forensic frameworks that exploit diffusion-model behavior—such as reconstruction dynamics or manifold alignment—remain limited.
@@ II-C Manifold-Based Forensics and Feature Design
Traditional image forensics relies on pixel- or frequency-level cues, including sensor noise and compression artifacts [13]. More recently, manifold-based reasoning has gained attention, suggesting that generative models learn implicit data manifolds where synthetic images lie on or near the manifold, while real images may deviate [15, 9]. Some approaches leverage reconstruction or projection errors in latent spaces; however, many require costly inversion procedures or assume knowledge of the generation model. As a result, generalization across unseen architectures and content types remains challenging.
II-D Research Gap and Motivation
Despite substantial progress in GAN detection and diffusion model analysis, key limitations persist. Existing detectors often lack cross-model generalization and robustness to real-world distortions. Many approaches rely on static image-level artifacts rather than analyzing dynamic reconstruction behavior. Furthermore, few methods treat diffusion models themselves as forensic probes for diverse content types beyond faces. These gaps motivate the development of dynamically informed features based on diffusion reconstruction trajectories that generalize across domains and generative architectures.
@@ III Methodology
Our detection approach analyzes reconstruction behavior under varying noise levels in diffusion models. For convenience, we refer to this behavior as diffusion snap-back.
@@ III-A Dataset
We used the AI vs. Human-Generated Images Dataset by Alessandra Sala, publicly available on Kaggle [14], as the primary dataset for our experiments. The dataset contains equal proportions of authentic (human-captured) and AI-generated images across multiple categories, including portraits, objects, and scenes. Each image is labeled as 0 (human-generated) or 1 (AI-generated), as shown in Table I. Representative samples are presented in Figure 1, where the top row shows real images and the bottom row shows AI-generated counterparts. Although the dataset is primarily face-centric, it includes diverse content such as portraits, landscapes, and objects, supporting evaluation across multiple domains. Authentic samples include photographs and original artwork, while synthetic samples are generated using Stable Diffusion v1.5.
@@ III-C Classification Pipeline
[…]
The overall detection pipeline is illustrated in Figure 2. The extracted features were used to train a lightweight logistic regression classifier with $\ell_{2}$ regularization. Prior to training, missing values were imputed using the median and features were standardized. Stratified five-fold cross-validation was conducted to assess generalization, yielding mean AUROC and AUPRC values. The optimal decision threshold $\theta^{*}$ was determined via Youden’s J-statistic, balancing sensitivity and specificity. To benchmark performance, a pixel-level baseline using 32×32 flattened image vectors achieved only 0.525 AUROC, underscoring the effectiveness of the proposed manifold features. On the full feature set, the snap-back model reached 0.993 cross-validation AUROC and 0.990 on a 35% holdout test split.
III-D Robustness Evaluation
To assess real-world robustness, we evaluated the model on six augmentation conditions applied to a balanced subset of 24 images (12 real, 12 AI-generated): JPEG-60, WebP-60, Gaussian blur (radius 1.2), additive noise ($\sigma=6.0$), screenshot resampling, and raw images. Snap-back features were recomputed for each variant and AUROC was measured. As shown in Table IV, performance remained stable under compression (83–87% AUROC), with moderate degradation under blur and screenshot distortions.
These results suggest that diffusion snap-back features retain discriminative power under common online perturbations.
III-E Implementation Details and Hyperparameters
All experiments were conducted in Python using PyTorch, the diffusers library, lpips, and scikit-learn. Images were first corrected for orientation, converted to RGB format, center-cropped, and resized to 512×512 to ensure consistent preprocessing.
Reconstructions were generated using the Stable Diffusion v1.5 img2img pipeline with a DDIM scheduler and 50 denoising steps. Four reconstruction strengths (0.15, 0.30, 0.60, and 0.90) were applied to simulate increasing levels of controlled perturbation. Mixed-precision inference was used on GPU to improve computational efficiency.
For each reconstruction strength, three perceptual similarity measures—LPIPS (AlexNet backbone), SSIM, and PSNR—were computed between the original and reconstructed images, resulting in 12 direct comparison features. To summarize reconstruction behavior across strengths, we derived three additional descriptors: the area under the LPIPS curve, the change in LPIPS between low and moderate strengths, and the first reconstruction level at which structural similarity dropped below a threshold of 0.80. Together, these formed a compact 15-dimensional feature representation.
@@ IV-A Primary Results
Table II summarizes the primary detection performance. Our diffusion snap-back feature pipeline achieves an AUROC of 0.993 on 5-fold cross-validation, and maintains 0.990 on a 35% holdout test set. The optimal decision threshold determined via Youden’s J-statistic is $\theta^{*}=0.914$, yielding balanced sensitivity and specificity.
The overall performance metrics are visualized in Fig. 5, which shows the ROC curve, calibration reliability plot, and confusion matrix at $\theta^{*}$. The model demonstrates high discriminative power and near-perfect calibration, with minimal false positives or negatives.
@@ IV-B Ablation Study
Table III presents the top-performing feature combinations. Ablation analysis identifies knee-step (the threshold at which SSIM drops below 0.8) as the single most discriminative feature. Combined with LPIPS at higher diffusion strengths and curve-level AUC summaries, the model approaches full-feature accuracy with minimal redundancy.
@@ IV-G Why Snap-Back Works
Diffusion models learn a denoising function $\nabla\log p_{\theta}(x)$ that approximates the data distribution through score matching. Images generated by the model are strongly aligned with this learned prior, whereas real-world photographs may occupy regions that are less precisely represented.
When noise is injected and reconstruction begins, this difference becomes apparent. Real images often diverge more sharply under high noise, resulting in larger LPIPS increases and faster SSIM degradation. In contrast, AI-generated images tend to recover more smoothly, maintaining structural consistency across reconstruction strengths.
This contrast in reconstruction dynamics reflects varying degrees of alignment with the learned denoising prior. As such, diffusion snap-back provides an interpretable signal for distinguishing synthetic images from authentic ones.
IV-H Deployment Considerations
The snap-back framework is lightweight and modular, requiring four img2img reconstructions followed by metric computation and a linear classifier, enabling transparent and efficient inference.
It can be deployed as an “upload-and-check” module within admission portals, recruitment systems, journalism workflows, or social media platforms to flag potentially synthetic submissions. Inference can be accelerated through batching and GPU parallelization, and the framework can be adapted to alternative diffusion backbones with minimal modification.
Overall, the separation between reconstruction dynamics and linear classification supports practical and interpretable real-world deployment.
@@ V Conclusion
This work introduced a diffusion-based forensic framework for distinguishing authentic photographs from AI-generated images by analyzing their reconstruction behavior under controlled noise strengths. Rather than relying on static pixel-level artifacts, we modeled multi-strength img2img reconstruction and extracted interpretable snap-back features aligned with the diffusion model’s learned denoising prior. The proposed approach demonstrates that reconstruction dynamics can serve as a complementary and principled signal for synthetic media detection.
Several limitations remain. Experiments were conducted using a single diffusion backbone, and broader cross-model validation is necessary to assess generalization. Hyperparameter exploration was limited, and more systematic tuning of reconstruction strengths and thresholds may further refine performance. Additionally, larger and more diverse datasets are required to better understand robustness across domains.
Future work will focus on cross-diffusion evaluation, expanded hyperparameter optimization, and scaling reconstruction analysis with improved computational efficiency. Extending snap-back analysis to video also represents a promising direction for advancing reliable and scalable synthetic media forensics.